Assignment:
Suppose that in a certain malarious area past experience indicates that the
probability of a person with a high fever will be positive for malaria is 0.7. Consider 3 randomly selected patients (with high fever) in that same area. 1) What is the probability that no patient will be positive for malaria?
2) What is the probability that exactly one patient will be positive for malaria?
3) What is the probability that exactly two of the patients will be positive for
malaria?
4) What is the probability that all patients will be positive for malaria?
5) Find the mean and the SD of the probability distribution given above.

Answer by Boreal(15235)

the mean is np=2.1 cases
variance is np(1-p)=0.63 cases^2
sd is sqrt (V)=0.79 cases
-
no positive patients is 0.3^3=0.027
exactly 1 positive is 3*0.7*0.3^2 since three ways that can happen=0.189
exactly 2 positives is 3*0.7^2*0.3=0.441
exactly 3 positives is 0.7^3=0.343
Those 4 possibilities add to 1.0